OpenStudy (anonymous):

Chase wants to factor x2 + 10x + 25 by grouping; however, Paige says it is a special product and can factor a different way. Using complete sentences, explain and demonstrate how both methods will result in the same factors.

OpenStudy (anonymous):

Well I'm pretty sure there is no immediate GCF

OpenStudy (mathmale):

Hint: a^2 + 2ab + b^2 = (a+b)(a+b). Where have you seen this before?

OpenStudy (anonymous):

I've seen it before but I forget what its for

OpenStudy (anonymous):

So I take it that we use that formula? Which numbers are 'a' and which is 'b'?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\( x^2 + 10x + 25 \)

ganeshie8 (ganeshie8):

\( x^2 + 2\times5x + 5^2\)

ganeshie8 (ganeshie8):

compare it wid \(a^2 + 2ab + b^2\)

OpenStudy (anonymous):

doesn't the facots you choose for 10x have to have a sum equal to 10x?

ganeshie8 (ganeshie8):

we will do that later. For now we're working it using the short cut method : 'special product'

ganeshie8 (ganeshie8):

\(x^2 + 2\times5x + 5^2\) compare it wid \(a^2 + 2ab + b^2\) \(a = ?\) \(b = ?\)

OpenStudy (anonymous):

a is 2 and b is 5?

ganeshie8 (ganeshie8):

nope. a = x, b = 5

ganeshie8 (ganeshie8):

simply plug a, b values in the formula :- \(a^2+2ab+b^2 = (a+b)(a+b)\)

ganeshie8 (ganeshie8):

\(x^2 + 2\times5x + 5^2 = (x+5)(x+5)\)

ganeshie8 (ganeshie8):

thats how we do it using first method : 'special product method'

OpenStudy (anonymous):

now we simplify it?

ganeshie8 (ganeshie8):

We're done.

ganeshie8 (ganeshie8):

question is asking us to factor given polynomial. we factored the polynomial. So we're done. ok

OpenStudy (anonymous):

Ok, lemme right it down real quick :)

OpenStudy (anonymous):

next we have to factor by grouping I think

ganeshie8 (ganeshie8):

you need to do it using second method : 'Grouping method' also

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

\(x^2 + 10x + 25 \) find two numbers such that :- sum = 10 product = 25

OpenStudy (anonymous):

5 and 5

OpenStudy (anonymous):

5 + 5 = 10 and 5 * 5 = 25 :D

ganeshie8 (ganeshie8):

yup ! so group it

ganeshie8 (ganeshie8):

\(x^2 + 10x + 25\) \(x^2 + 5x + 5x + 25\)

ganeshie8 (ganeshie8):

factor the GCF from first two terms factor the GCF from last two terms

OpenStudy (anonymous):

x^2 + 5x + 5x + 25 shouldn't only 1 5 have an x? 5x + 5x = 10x^2 right?

ganeshie8 (ganeshie8):

5apples + 5apples = 10 apples

ganeshie8 (ganeshie8):

5x + 5x = 10x

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

next we can regroup right?

ganeshie8 (ganeshie8):

\(x^2 + 10x + 25\) \(x^2 + 5x + 5x + 25\) \(x(x+5) + 5(x+5)\)

OpenStudy (anonymous):

then we make the answer into (x + 5)(x + 5)?

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

then we done? :D

OpenStudy (anonymous):

yay we're done :D

OpenStudy (anonymous):

both answers are identical :)

ganeshie8 (ganeshie8):

yes we're done :)